### v1

# Ergodicity, KAM, FPUT

Abstract
[…]
1 Ergodicity
The Hamiltonian of the systems of $N$ oscillators $r_{x}=(p_{x},r_{x}),x=1,...,N$ , considered by FPUT was, ($N=64,r_{0}=r_{N}=0$):
$\sum^{N-1}_{x=1}\frac{p_{x}^{2}}{2m}+\sum_{x=1}^{N-1}\frac{1}{2}(r_{x+1}-r_{x})^{2}+\frac{\alpha}{3}(r_{x+1}-r_{x})^{3}+\frac{\beta}{4}(r_{x+1}-r_{x})^{4}$ (1.1)
representing a linear chain of oscillators, which can be generalized to a $d$-dimensional lattice with $r,x$ as $d$-dimensional vectors, for instance:
$\sum_{{\bf x}\in Z_{N}}\frac{1}{2}{{\bf p}^{2}_{\bf x}}+\sum_{{\bf x}\in Z_{N}}\sum_{i=1}^{d}\frac{1}{2}({\bf r}({\bf x}+{\bf e}_{i})-{\bf r}_{\bf x})^{2}+\frac{\beta}{2}({\bf r}_{(}{\bf x}+{\bf e}_{i})-{\bf r}_{\bf x})^{4}$ (1.2)
where ${\bf e}_{i},i=1,\ldots,d$ is an ortonormal basis and $Z_{N}=(0,\ldots,N-1)^{d}$ is the cube with side $N$ in the lattice $Z^{d}$, and ${\bf r}_{\bf x}=0$ if any coorinate of $\bf x$} is 0 or $N$.
Boltzmann’s ergodic hypothesis, [6], formally stated an hypothesis, which had been perhaps earlier conceived (for instance by Descartes, [7, p.4]) and which had been accepted by Maxwell, [8], supposes that the point representing the system istantaneous state wonders visiting all points that can possibly be visited (i.e. compatible with the constraints).
The hypothesis, leaving aside the obvious exceptions, and its implication that all points covered in time a closed orbit allowed Boltzmann to discover a mechanical interpretation of the Second Law as a manfestation of the Maupertuis’ least action principle via the following extraordinary argument.
(i) An equilibrium state of a system is idetified with a periodic motion $t\to x(t)$ developing under the action of internal forces with potential energy $V(x(t))$ and of external forces with potential $V_{ext}(x(t))$; its kinetic energy is $K(x(t))$, and the action of $x$ is defined, if its period is $i$ (adopting here the original notation for the period), by
${\cal A}(x)=\int_{0}^{i}\big{(}\frac{m}{2}\dot{x}(t)^{2}-V(x(t))\big{)}\,dt$ (1.3)
(ii) Properties of a small periodic variations $\delta x(t)$ of a state, i.e. properties of an equilibrium state close to $x$, represented as:
$\delta x(t)=x^{\prime}(\frac{i^{\prime}}{i}t)-x(t){\buildrel def\over{=}}x^{\prime}(i^{\prime}\varphi)-x(i\varphi)\hfill$ (1.4)
where $\varphi\in[0,1]$ is the phase, as introduced by Clausius, [9]. The role of $\varphi$ is simply to establish a correspondence between points on the initial trajectory $x$ and on the varied one $x^{\prime}$: it is manifestly an arbitrary correspondence (which could be defined differently without affecting the final result) convenient to follow the algebraic steps. [^1] It should be noted that Boltzmann does not care to introduce the phase and this makes his computations difficult to follow (in the sense that once realized which the final result is, it may be easier to reproduce it rather than to follow his calculations).
(iii) Following Clausius version, [9], the heat theorem is deduced from a few identities about the average values ${\overline{F}}=i^{-1}\int_{0}^{i}F(x(t))dt$ for generic observables $F$ and ${\overline{K}}=i^{-1}\int_{0}^{i}\frac{1}{2}\dot{x}(t)^{2}dt$ for the average of the kinetic energy.
Let $x(t)=\xi(\frac{t}{i})$ with $\xi(\varphi)$ 1-periodic; it is ${\overline{F}}\equiv\int_{0}^{1}F(\xi(\varphi))d\varphi$, and ${\overline{K}}=\frac{1}{2}\int_{0}^{1}\frac{\xi^{\prime}(\varphi)^{2}}{i^{2}}d\varphi$. Changing $x$ by $\delta x$ means to change $\xi$ by $\delta\xi$ and $i$ by $\delta i$. So variation of an average is the sum of the variation due to $\delta\xi$ and that due to $\delta i$: namely $\delta{\overline{F}}=\delta_{\xi}{{\overline{F}}}+\delta_{i}{\overline{F}}$. Therefore
${\delta{\overline{V}}\equiv\delta_{\xi}{{\overline{V}}}},\qquad\delta{\overline{V}_{ext}}=\delta_{\xi}{{\overline{V}_{ext}}},\qquad{\delta{\overline{K}}=-2\frac{\delta i}{i}{\overline{K}}+\delta_{\xi}{\overline{K}}}$ (1.5)
The variation $\delta({\overline{K}}+{\overline{V}}+{\overline{V}_{ext}}){\buildrel def\over{=}}dU+dL$, with $U{\buildrel def\over{=}}{\overline{K}}+{\overline{V}}$ and $dL{\buildrel def\over{=}}d{\overline{V}_{ext}}$, has a natural interpretation of heat $\delta Q$ in/out in the process $x\to x^{\prime}$.  [^2] Because $\delta({\overline{K}}+{\overline{V}}){\buildrel def\over{=}}dU$ and $\delta{\overline{V}_{ext}}{\buildrel def\over{=}}\delta L$ are the variation of the internal energy $U$ and $dL$ is the work that the system performs: i.e. $\delta Q=\delta U+\delta L$.
Then using the above Eq. (1.5): $\delta_{\xi}{\overline{K}}=\delta{\overline{K}}+2\frac{\delta i}{i}{\overline{K}}$
${{}\delta Q\equiv{\delta({\overline{K}}+{\overline{V}}+{\overline{V}_{ext}})=-2\frac{\delta i}{i}{\overline{K}}+\delta_{\xi}({{\overline{K}}}+{\overline{V}}+{\overline{V}_{ext}})}}$
${{}={\bf-}2\frac{\delta i}{i}{\overline{K}}+{2\delta_{\xi}{{\overline{K}}}}+{\mbox{\boldmath$\delta$}}_{\xi}{\bf({-{{\overline{K}}}}+{\overline{V}}+{\overline{V}_{ext}})}\equiv{\bf+}2\frac{\delta i}{i}{\overline{K}}+2\delta{{\overline{K}}}+{\bf 0}}$
where the last $\bf 0$} is due to Maupertuis’ principle, as the motion follows the equations of motion (at fixed $i$). Therefore, (Heat theorem):
${\frac{\delta Q}{{\overline{K}}}}=+2\frac{\delta i}{i}+2\frac{\delta{\overline{K}}}{{\overline{K}}}={2\delta\log(i{\overline{K}})}{\buildrel def\over{=}}{\bf\delta S}$ (1.7)
is an exact differential. In reality it is somewhat strange that both Boltzmann and Clausius call the last equation a “generalization of the action principle”: the latter principle uniquely determines a motion, i.e. it determines its equations. The equation instead does not determine a motion but it only establishes a relation between the variation of average kinetic and potential energies of close periodic motions under the assumption that they satisfy the equations of motion; and it does not establish a variational property, see [10, Sec.I.4].
The above analysis was successive to the ergodic hypothesis and a complement to it: after achieving the construction of the microcanonical ensemble, [1], as recognized by Maxwell in one of his last works,[8], and later by Gibbs, [11] (who refers to a later publication, [12], curiously quoting it under the title of the first section) and soon Boltazmann wanted to provide simple exapmles of ergodic systems.
He examined the system of one particle, in dimension $2$, subject to a central force of potential $\frac{\alpha}{|{\bf x}|}$, to a potential $\frac{\beta}{|{\bf x}|^{2}}$ and to a flat repelling line. For $\beta=0$ the motion is integrable and Boltzmann tried to prove that as soon as $\beta>0$ the system would be ergodic. This is not right (as conjectured in [13, 14] and proved in [15]) the example fails as a consequence of the KAM theorem, discovered in the early ’950s by Kolmogorov, [4]. Nevertheless Boltzmann’s paper was probably known to Fermi, who in his early youth attempted to provide a proof of the ergodic hypothesis,[16] (overlooking a similar difficulty: i.e. the possibility of the existence of non smooth constants of motion); hence when (relatively) large computers became available the FPUT project was attempted.
Existence of non smooth constants of motion rests, ultimately, upon realizing that the phase space may have a large volume covered by smooth invariant surfaces of low dimension: however proving their existence requires a multi scale analysis: by examining the motions with ever increasing precision, invariant curves can be separated from the nearby motions (which are likely to undergo chaotic motions).
2 KAM
In spite of its interest the first attempt to establish a rigorous connection between the problem of existence of non smooth constant of motions and the ergodicity of the FPUT chain had to wait 1971, [17], treating “only” $N$ particles chains with $N$ either prime or a power of $2$ (hence including at least the cases studied in FPUT work N=16,32,64N=16,32,64!) and a complete proof, including the “other $N$-s” (NN not a power of $2$ or not a prime), dates 2006, [18]; both, although conceptually very different, relied on the KAM theorem, which is discussed below as an application of the renormalization group method.
Consider the Hamiltonian:
$H_{0}({\bf A},{\mbox{\boldmath$\alpha$}})=\frac{1}{2}({\bf A}\cdot J_{0}{\bf A})+{\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}+f_{0}({\bf A},{\mbox{\boldmath$\alpha$}})\equiv h_{0}+f_{0}$ (2.1)
real analytic for $({\bf A},{\mbox{\boldmath$\alpha$}})\in({\mathcal{D}}_{\rho}\times{\mathcal{T}}^{N})$, $N\geq 2$, with: ${\mathcal{D}}_{\rho}=\{{\bf A}\in R^{N},|A_{j}|<\rho\}$, ${\mathcal{T}}^{N}$ = $N$-dimensional torus $0,2\pi]^{N}$, identified with unit polydisk, $\{{\bf z}|\,z_{j}=e^{i\alpha_{j}},j=1,\ldots,N\}$, ${\mbox{\boldmath$\omega$}}_{0}\in R^{N}$ and $J_{0}$ is a $N\times N$ matrix, positive for simplicity, with eigenvalues in $J_{0,-},J_{0,+}]$.
The Hamiltonian is supposed holomorphic in the complex region ${\mathcal{C}}_{\rho_{0},\kappa_{0}}=\{({{\bf A}},{\bf z})||A_{j}|\leq\rho_{0},e^{-\kappa_{0}}\leq|z_{j}|\leq e^{\kappa_{0}}\}$; the label 0 is attached, since the beginning, because $H_{n},f_{n},\rho_{n},\kappa_{n}$ will arise later, with $n=1,2,\ldots$; the perturbation $f_{0}$ size will be measured by $\varepsilon_{0}$ as defined here:
${{\mathcal{C}}_{\rho_{0},\kappa_{0}}}$ ${{}{\buildrel def\over{=}}\{({\bf A},{\bf z})||A_{j}|\,\leq\,\rho_{0},\ e^{-\kappa_{0}}\leq|e^{i\alpha_{j}}|\leq e^{\kappa_{0}},j=1,\ldots,N\}\subset{\mathcal{C}}^{2N}}$
[…]
${||f}$ ${{}||_{\rho_{0},\kappa_{0}}{\buildrel def\over{=}}\max_{{\mathcal{C}}_{\rho_{0},\kappa_{0}}}|f({\bf A},{\bf z})|,\qquad\forall f\ {\rm holomorphic\ in}\ {\mathcal{C}}_{\rho_{0},\kappa_{0}}}$
with $\rho_{0}>0,1\geq\kappa_{0}>0$, $z_{j}\equiv e^{i\alpha_{j}}$; generally ${\mathcal{C}}_{\rho,\kappa}({\bf{\overline{A}}})$ will denote a ’polydisk’ centered at $\bf{\overline{A}}$}, i.e. defined as in Eq.(2.2) with $|A_{j}-{\overline{A}}_{j}|\leq\rho$ replacing $|A_{j}|\leq\rho$ and $\kappa$ replacing $\kappa_{0}$; polydisks centered at the “origin” ${\overline{{{\bf A}}}}=0$ will be simply denoted ${\mathcal{C}}_{\rho,\kappa}$ and called “centered polydisks”, while $C_{\rho}({\overline{{{\bf A}}}})$ is the $\{{{\bf A}}||A_{i}-{\overline{A}}_{i}|<\rho\}$. Set $|{\bf A}|=\max|A_{j}|,|{\bf z}|=\max|z_{j}|,\ \forall{\bf A},{\bf z}\in{\mathcal{C}}^{N}$.
For simplicity functions of $\alpha$ will be implicitly regarded as functions of $z_{j}=e^{i\alpha_{j}}$ and their arguments will be written as $\bf z$} or $\alpha$, as convenient.
The idea is to focus attention on the center of ${\mathcal{C}}_{\rho_{0},\kappa_{0}}$ where, if $\varepsilon_{0}=0$, a motion (“free motion”: $({\bf A},{\mbox{\boldmath$\alpha$}})\to({\bf A},{\mbox{\boldmath$\alpha$}}+{\mbox{\boldmath$\omega$}}_{0}t)$) takes place which is quasi periodic “with spectrum” ${\mbox{\boldmath$\omega$}}_{0}$. This is done by studying motions in a small polydisk ${\mathcal{C}}_{\widetilde{\rho},\widetilde{\kappa}}({\bf a})\subset{\mathcal{C}}_{\rho_{0},\kappa_{0}}$, eccentric if ${\bf a}\neq{\bf 0}$.
Thus motions starting in ${\mathcal{C}}(\widetilde{\rho},\widetilde{\kappa})({\bf a})\subset{\mathcal{C}}(\rho_{0},\kappa_{0})$ near the center of ${\mathcal{C}}(\rho_{0},\kappa_{0})$ can be studied as “through a microscope”: in the good cases (i.e. under suitable assumption on the initial parameters $J_{0},{\mbox{\boldmath$\omega$}}_{0}$ and $f_{0}$) the Hamiltonian will turn out to be expressible (after a further coordinate change, to turn ${\mathcal{C}}(\widetilde{\rho},\widetilde{\kappa})({\bf a})$ into ${\mathcal{C}}(\rho_{1},\kappa_{1})$, centered at 0) in a form substantially closer to that of a quasi periodic oscillator (described by its “normal” Hamiltonian ${\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}$, or ${\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}+\frac{1}{2}{{\bf A}}\cdot J_{0}{{\bf A}}$, in the variables ${\bf A},{\mbox{\boldmath$\alpha$}}$).
Iterating the process the Hamiltonian changes but, remaining analytic in a polydisk $C_{\rho_{n},\kappa_{n}}$, converges to that of a harmonic oscillator: the interpretation will be that, looking very carefully in the vicinity of the “unperturbed” torus ${\mathcal{T}}_{{\mbox{\boldmath$\omega$}}_{0}}=\{{\bf A}={\bf 0},{\mbox{\boldmath$\alpha$}}\in[0,2\pi]^{N}\}$, the perturbed Hamiltonian exhibits with increasing precision a harmonic motion with spectrum ${\mbox{\boldmath$\omega$}}_{0}$.
The result is the KAM theorem in a form not only reminiscent of the methods called “renormalization group”, RG, in quantum field theory but just a realization of them, [19]. Existence of the invariant torus can be seen as the existence of a trivial fixed point (a harmonic oscillator) of a single canonical coordinate change together with a time rescaling.
The “spectrum” ${\mbox{\boldmath$\omega$}}_{0}$ will be supposed “diophantine”, i.e. for some $C>0$ it is, denoting $\mathcal{Z}^{N}$ the lattice of the integers, for all ${\bf 0}\neq{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N}$:
$|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|\,\geq\,C\,|{\mbox{\boldmath$\nu$}}|^{-N},\qquad|{\mbox{\boldmath$\nu$}}|\equiv\sum_{i=1}^{N}|\nu_{i}|>0$ (2.3)
The inequality Eq.2.3 will be repeatedly used to define canonical transformations $({{\bf A}},{\mbox{\boldmath$\alpha$}})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ with generating functions $\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})+({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\alpha$}}$:
${\bf A}={\bf A}^{\prime}+{\bf a}+{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}),\quad{\mbox{\boldmath$\alpha$}}^{\prime}={\mbox{\boldmath$\alpha$}}+{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})$ (2.4)
with $\Phi$ chosen so that in the new coordinates $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ the perturbation is ’weaker’; at the price that the new coordinates will cover a (much) smaller domain, inside ${\mathcal{C}}(\rho_{0},\kappa_{0})$.
[…]
Suppose that a holomorphic $\Phi({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ defines, via Eq.(2.4), in a domain $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})\in{\mathcal{C}}_{\widetilde{\rho},\widetilde{\kappa}}$, a symplectic map $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})\to({\bf A},{\mbox{\boldmath$\alpha$}})$ with image $\subset{\mathcal{C}}_{\rho_{0},\kappa_{0}}$.
Choose $\bf a$} as solution of the implicit equation ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ and insert Eq.(2.4) into $H_{0}$.
After easy&patient algebra, the Hamiltonian becomes,  444 $\alpha$ has to be imagined as a function of $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$, solution of the second implicit equation in Eq.2.4.  up to a constant, the sum of the four lines (that will be referred as $F_{j},j=0,\ldots,4$):
${(0):\quad}$ ${{}H^{\prime}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=\frac{1}{2}({\bf A}^{\prime}\cdot J_{0}{\bf A}^{\prime})+{\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}^{\prime}}$
${(1):\quad}$ ${{}+{\overline{f}}_{0}({\bf A}^{\prime}+{\bf a})-{\overline{f}}_{0}({\bf a})-{\mbox{\boldmath$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})\cdot{\bf A}^{\prime}}$
${(2):\quad}$ ${{}+\Big{(}f_{0}({\bf A}^{\prime}+{\bf a}+{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi,{\mbox{\boldmath$\alpha$}})-f_{0}({\bf A}^{\prime}+{\bf a},{\mbox{\boldmath$\alpha$}})\Big{)}+\frac{1}{2}{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi\cdot J_{0}{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi}$
${(3):\quad}$ ${{}+({\mbox{\boldmath$\omega$}}_{0}+J_{0}({\bf A}^{\prime}+{\bf a}))\cdot{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\alpha$}}}\Phi+f_{0}({\bf A}^{\prime}+{\bf a},{\mbox{\boldmath$\alpha$}})-{\overline{f}}_{{\bf 0}}({\bf A}^{\prime}+{\bf a})}$
where a few terms have been added or subtracted (including free addition or subtraction of constants, like ${\mbox{\boldmath$\omega$}}_{0}\cdot{\bf a}$, $\frac{1}{2}J_{o}{\bf a}\cdot{\bf a}$ and ${\overline{f}}_{0}({\bf a})$): so that each line $F_{j},j>0$, can formally be estimated to have size $\sim O(\varepsilon_{0}^{2})$ in a smaller polydisk, while the line $(0)$ is the new unperturbed Hamiltonian (in the $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ variables 4 ).  [^5] Heuristically: line (1) is a sum of quantities of order $\varepsilon_{0}$ but it is of second order in $\bf A$$^{\prime}$ and could be made of $O(\varepsilon_{0}^{2})$ by suitably restricting size of $\bf A$$^{\prime}$: remark that ${\overline{f}}_{0}({\bf a})$ is a constant added freely but the last term is there because of the choice ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})$. Line (2) can be made of $O(\varepsilon_{0}^{2})$ by choosing $\Phi$ of $O(\varepsilon_{0})$. Line (3) could even be set =0=0 and invites taking it a definition of $\Phi$ (if possible).
As familiar in pertubation theory this may suggest to define:
${{}{\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})}$
${{}\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})=-\sum_{{\bf 0}\neq{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N}}\frac{f_{0,{\mbox{\boldmath$\nu$}}}({\bf A}^{\prime}+{\bf a})}{i({\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}+(({\bf A}^{\prime}+{\bf a})\cdot J_{0}{\mbox{\boldmath$\nu$}}))}e^{i{\mbox{\boldmath$\nu$}}\cdot{\mbox{\boldmath$\alpha$}}},}$
Existence of $\bf a$} inside $C_{\rho_{0}}$ can be easily established via dimensional bounds.
  • In general an example of dimensional bound is a bound of a holomorphic function $f(\zeta)$ and of its derivatives at a point $\zeta$ in terms of its maximum $|f|$ in the definition domain and of the distance $d(\zeta)$ of $\zeta$ to the boundary: the $n^{th}$-derivative at $\zeta$ is bounded by $n!(\max|f|)\,d(\zeta)^{-n}$. Likewise a function $g(\zeta)$ holomorphic in an annulus bounded by the circles $\zeta=e^{i\alpha\pm\kappa},\alpha\in[0,2\pi]$, will have Fourier’s coefficients bounded by its maximum $\varepsilon$ in the annulus as $|g_{{\mbox{\boldmath$\nu$}}}|\leq\varepsilon\exp(-\kappa\,|{\mbox{\boldmath$\nu$}}|)$. The mentioned bounds will be called dimensional bounds.
  • In conclusion, if $\varepsilon=\max_{{\mathcal{C}}(\rho,\kappa)}|{\mbox{\boldmath$\partial$}}_{{\bf A}}F({{\bf A}},{\mbox{\boldmath$\alpha$}})|+\max_{{\mathcal{C}}(\rho,\kappa)}\frac{1}{\rho}|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}F({{\bf A}},{\mbox{\boldmath$\alpha$}})|$ and $F$ is holomorphic in ${\mathcal{C}}(\rho,\kappa)$, a dimensional bound on $F({{\bf A}},{\mbox{\boldmath$\alpha$}})$ for $({{\bf A}},{\mbox{\boldmath$\alpha$}})\in{\mathcal{C}}((1-\lambda)\rho,\kappa)$, $\lambda<1$, is, for ${\mbox{\boldmath$\nu$}}\neq 0$:
    $|{\mbox{\boldmath$\partial$}}_{{\bf A}}^{\bf n}{\mbox{\boldmath$\partial$}}_{{\bf A}}F({{\bf A}},{\mbox{\boldmath$\alpha$}})|<{\bf n}!\,(\lambda\rho)^{-|{\bf n}|}\,\varepsilon,\qquad|{\mbox{\boldmath$\nu$}}||F_{\mbox{\boldmath$\nu$}}({{\bf A}},{\mbox{\boldmath$\alpha$}})|<e^{-\kappa\,|{\mbox{\boldmath$\nu$}}|}\,\rho\,\varepsilon$ (2.7)
    where ${\bf n}\in\mathcal{Z}_{+}^{N},\,|{\bf n}|=\sum n_{i},\,{\bf n}!=\prod n_{i}!,\,{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N},\,|{\mbox{\boldmath$\nu$}}|=\sum_{i}|\nu_{i}|$.  [^6] Remark that to estimate $\bf A$}}-derivatives the $\rho$ has to be reduced by a factor $1-\lambda$ while to estimate $\alpha$-derivatives $\kappa_{0}$ can be left unchanged, as such $m$-th $\alpha$-derivatives modify the second of Eq.2.7 by a factor $|{\mbox{\boldmath$\nu$}}|^{m}$.
  • In the following, dimensional bounds will be used repeatedly: when a bound on a function holomorphic in ${\mathcal{C}}(\rho,\kappa)$ is needed, it will be obtained dimensionally but restricted to a smaller polydisk ${\mathcal{C}}(\lambda\rho,\kappa-\delta)$, for $\lambda<1,\delta<\kappa$ (conveniently fixed each time, without attention to optimize the choice to obtain better bounds).
  • As a first example, a condition for the existence of the shift $\bf a$}, in the first of Eq.(2.6), is $\varepsilon_{0}$ small: if $J_{0,-}^{-1}\varepsilon_{0}\rho_{0}^{-1}<\frac{1}{\gamma_{0}}$ and $\gamma_{0}$ is large enough, the equation ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}({\bf a})$ has a solution $\bf a$} with $|{\bf a}|<J_{0,-}^{-1}\varepsilon_{0}<\frac{1}{\gamma_{0}}\rho_{0}$: see Appendix A  where the choice $\gamma_{0}=3N$ is proposed.
However the $\Phi$ in Eq.(2.6) does not make sense (in general); instead it will “be approximated”, and properly defined, by $\Phi_{0}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})$ as:
${{}-\kern-8.0pt\sum_{{\bf 0}\neq{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N}}\kern-5.0pt\frac{f_{0,{\mbox{\boldmath$\nu$}}}({\bf A}^{\prime}+{\bf a})e^{i{\mbox{\boldmath$\alpha$}}\cdot{\mbox{\boldmath$\nu$}}}}{i{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}}{\Big{(}1-\frac{J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\nu$}}}{{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}}+\Big{(}\frac{J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\nu$}}}{{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}}\Big{)}^{2}\Big{)}}}$
From $\varepsilon_{0}=\|{\partial}_{{{\bf A}}^{\prime}}f_{0}\|_{\rho_{0},\kappa_{0}}+\frac{1}{\rho_{0}}\|{\partial}_{{\mbox{\boldmath$\alpha$}}}f_{0}\|_{\rho_{0},\kappa_{0}}$ the dimensional inequalities (Eq.2.7):
$|{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}f_{0,{\mbox{\boldmath$\nu$}}}({{\bf A}}^{\prime}+{\bf a})|\leq\varepsilon_{0}e^{-\kappa_{0}|{\mbox{\boldmath$\nu$}}|},\qquad\rho_{0}|{\mbox{\boldmath$\nu$}}||f_{0,{\mbox{\boldmath$\nu$}}}({{\bf A}}^{\prime}+{\bf a})|<\varepsilon_{0}e^{-\kappa_{0}\cdot|{\mbox{\boldmath$\nu$}}|}$ (2.9)
hold if ${{\bf A}}^{\prime}+{\bf a}\in C(\rho_{0})$; and allow to estimate $\Phi_{0}$ and to define properly the canonical map in Eq.2.4 as a map from ${\mathcal{C}}(\frac{1}{4}\rho_{0},\kappa_{0}-2\delta_{0})$ to ${\mathcal{C}}(\rho_{0}-\frac{1}{2}\rho_{0},\kappa_{0}-\delta_{0})$ for some analyticity loss parameter $0<\delta_{0}<\frac{1}{4}\kappa_{0}$.
This is done via a few dimensional bounds obtained by successive reduction of analiticity domains ${\mathcal{C}}(\rho,\kappa)$ to ${\mathcal{C}}(\rho^{\prime},\kappa^{\prime})$ conveniently chosen.  [^7] Warning: every time a new estimate is proposed, the dimensionless constants involved in it will typically be $\gamma_{j},c_{j}$ with a new label $j$ even when equal to constants with label $j^{\prime}<j$.
If $\gamma_{0}=3N$, as above, and $C$ is the Diophantine constant, Eq.(2.3), the
$J_{0,-}^{-1}\varepsilon_{0}<\frac{1}{\gamma_{0}}\rho_{0},\qquad J_{0,+}\rho_{0}C^{-1}<1$ (2.10)
and the second of Eq.(2.9) imply a simple bound on $\|\Phi_{0}\|_{\frac{2}{3}\rho_{0},\kappa_{0}}$  [^8] As ${{\bf A}}^{\prime}\in C(\frac{2}{3}\rho_{0})$, implies ${{\bf A}}^{\prime}+{\bf a}\in C(\frac{5}{6}\rho_{0})$: remark $|{\bf a}|<\frac{1}{3N}\rho_{0}<\frac{1}{6}\rho_{0}$ for $N\geq 2$. . Therefore
${{}\|\Phi_{0}\|_{\frac{2}{3}\rho_{0},\kappa_{0}}\leq\varepsilon_{0}\rho_{0}\kern-5.69054pt\sum_{{\mbox{\boldmath$\nu$}}\neq{\bf 0}}\frac{|{\mbox{\boldmath$\nu$}}|e^{-\kappa_{0}|{\mbox{\boldmath$\nu$}}|}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{(}1\kern-5.69054pt+\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}+\kern-5.69054pt\Big{(}\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{)}^{2}\Big{)}\leq\gamma_{1}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-c_{1}}}$
which yields
[…]
${{}\leq\widetilde{\gamma}\varepsilon_{0}\sum_{{\mbox{\boldmath$\nu$}}\neq{\bf 0}}\frac{|{\mbox{\boldmath$\nu$}}|^{2}e^{-\kappa_{0}|{\mbox{\boldmath$\nu$}}|}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{(}1+\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}+\Big{(}\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{)}^{2}\Big{)}\leq\widehat{\gamma}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}}$
with $\widetilde{\gamma},\widehat{\gamma},\widehat{c}$ dimensionless constants.  [^9] E.g. $\widehat{c}=4N+2$, $\widehat{g}=\widetilde{\gamma}\sum_{\mbox{\boldmath$\nu$}}|{\mbox{\boldmath$\nu$}}|^{4N+2}e^{-\kappa_{0}\,|{\mbox{\boldmath$\nu$}}|}$, $\widetilde{\gamma}=7$. Remark that the bound on the second term in the first line of Eq.2.12 does not involve a change in $\kappa_{0}$, unlike the first term (where $\rho$ is reduced).
[…]
The ’new’ perturbation will be, by Eq.(2.5), $f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ if expressed in the mixed variables $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$. Hence it will eventually be necessary to express $\alpha$ via $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ by inverting the second in Eq.2.4 as ${\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$: so that the perturbation acting in $H_{1}$ will be, properly, $f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))$.
To solve for the implicit function $\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ in the second of Eq.2.4 , an argument can be used similar to the one for the solution of ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ in Appendix A . In Appendix B the similar analysis is adopted to construct $\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ and a proof that it can be defined on ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ subject to the bound:
$\|\Delta\|_{\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0}}<{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-{\overline{c}}}$ (2.13)
for dimensionless ${\overline{\gamma}},{\overline{c}}$, and for $\delta_{0}>0$ to be determined and depending on $\varepsilon_{0},\kappa_{0}$. However it is convenient to postpone the study of $\Delta$ and to obtain first bounds on $f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ in a domain ${\mathcal{C}}(\widetilde{\rho},\kappa_{0})$. The conditions for the definitions of $F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, i.e. for the definition of the generating function $\Phi_{0}$ to be defined in the mixed variables $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, are summarized in Eq.2.10
Under such conditions the $F_{j}$ composing $f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, in the lines labeled $(j),j>0$ in Eq.(2.5), can be easily bounded dimensionally as follows.
  • (a) $F_{0}$ is $\frac{1}{2}(J_{0}{{\bf A}}^{\prime}\cdot{{\bf A}}^{\prime})+{\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}^{\prime}$ for ${{\bf A}}^{\prime}\in C(\frac{1}{2}\rho_{0})$.
[…]
    $\|F_{1}\|_{\widetilde{\rho},\kappa_{0}}\leq\widetilde{\gamma}_{1}\varepsilon_{0}\frac{\widetilde{\rho}^{2}}{\rho_{0}}\qquad{\bf A}^{\prime}\in{\mathcal{C}}(\widetilde{\rho},\kappa_{0}),\quad\widetilde{\rho}\leq\frac{1}{2}\rho_{0}$ (2.14)
    with $\widetilde{\gamma}_{1}$ dimensionless. Hence $\|{\partial}_{{{\bf A}}^{\prime}}F_{1}\|_{\widetilde{\rho}}\leq\gamma_{1}\varepsilon_{0}\frac{\widetilde{\rho}}{\rho_{0}}$, $\widetilde{\rho}<\frac{1}{4}\rho_{0},\gamma_{1}=4\widetilde{\gamma}_{1}$.
  • (c) For $\widetilde{\rho}\leq\frac{1}{4}\rho_{0}$ by Eq. 2.52.12  it is $|F_{2}|_{\frac{1}{2}\rho_{0},\kappa_{0}}<\varepsilon_{0}4\rho_{0}\gamma_{2}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-c_{2}}$ in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0})$. Hence, for $\gamma_{3},c_{3}$ dimensionless:
    ${{}|{\partial}_{{{\bf A}}^{\prime}}F_{2}|_{\widetilde{\rho},\kappa_{0}}+\frac{1}{\widetilde{\rho}}|{\partial}_{\mbox{\boldmath$\alpha$}}F_{2}|_{\widetilde{\rho},\kappa_{0}}\leq\gamma_{3}\varepsilon_{0}\frac{\rho_{0}}{\widetilde{\rho}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-c_{3}}}$
    on the two contributions to $F_{2}$ (remark that $\rho_{0}-\frac{1}{4}\rho_{0}>\frac{1}{4}\rho_{0}$ and that only the first is responsible of the factor $\frac{\rho_{0}}{\widetilde{\rho}}$).  [^10] Remark: the domain in $|{{\bf A}}^{\prime}|$ has been reduced to $\frac{1}{4}\rho_{0}$ to permit the dimensional estimate of the derivative of the first of the two terms in $F_{2}$, as is done below also to bound $F_{3}$.
  • (d) Finally $F_{3}$, with $\Phi_{0}$ defined in Eq.(2.8), does not vanish, but an occurring algebraic cancellation simplifies it, under the conditions in Eq.2.10 and, using Eq. 2.9 , as:
    ${{}F_{3}({\bf A}^{\prime},{\bf z}){\buildrel def\over{=}}({\mbox{\boldmath$\omega$}}_{0}\kern-2.84526pt+\kern-2.84526ptJ_{0}({\bf A}^{\prime}\kern-2.84526pt+\kern-2.84526pt{\bf a})\cdot{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi_{0}\kern-2.84526pt+\kern-2.84526ptf_{0}({\bf A}^{\prime}\kern-2.84526pt+\kern-2.84526pt{\bf a},{\mbox{\boldmath$\alpha$}})\kern-2.84526pt-\kern-2.84526pt{\overline{f}}_{0}({\bf A}^{\prime}+{\bf a}))}$
    ${{}=-\sum_{{\bf 0}\neq{\mbox{\boldmath$\nu$}}}f_{0,{\mbox{\boldmath$\nu$}}}({\bf A}^{\prime}+{\bf a})\frac{(J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\nu$}})^{3}}{({\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}})^{3}}e^{i{\mbox{\boldmath$\alpha$}}\cdot{\mbox{\boldmath$\nu$}}}\leq\widetilde{\gamma}\frac{\varepsilon_{0}}{\rho_{0}}\widetilde{\rho}^{3}\kappa_{0}^{-c}}$
    and can be bounded in ${\mathcal{C}}_{\widetilde{\rho},\kappa_{0}}$ by $\widetilde{\gamma}\frac{\varepsilon_{0}}{\rho_{0}}\widetilde{\rho}^{3}\kappa_{0}^{-c}$ for suitable dimendionless $\widetilde{\gamma},c$, and in the variables $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\in{\mathcal{C}}_{\widetilde{\rho},\kappa_{0}}$ as in c) above:
    ${{}|{\partial}_{{{\bf A}}^{\prime}}F_{3}|_{\widetilde{\rho},\kappa_{0}}+\frac{1}{\widetilde{\rho}}|{\partial}_{\mbox{\boldmath$\alpha$}}F_{3}|_{\widetilde{\rho},\kappa_{0}}\leq\gamma_{4}\varepsilon_{0}\Big{(}\frac{\widetilde{\rho}}{\rho_{0}}\Big{)}^{2}\kappa_{0}^{-c_{4}}}$
    where $\gamma_{4},c_{4}$ are dimensionless constants and $\widetilde{\rho}\leq\frac{1}{4}\rho_{0}$.
Therefore if $\widetilde{\rho}\leq\frac{1}{4}\rho_{0},\widetilde{k}\leq\kappa_{0}$ and if Eq.2.10 holds:
${{}J_{1}=J_{0},\qquad|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}F_{1}({\bf A}^{\prime})|\leq{\overline{\gamma}}\varepsilon_{0}\Big{(}\frac{\widetilde{\rho}}{\rho_{0}}\Big{)},}$
${{}|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}F_{2}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|+\frac{1}{\widetilde{\rho}}|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}F_{2}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|\leq{\overline{\gamma}}\varepsilon_{0}\frac{\varepsilon_{0}}{C}\frac{\rho_{0}}{\widetilde{\rho}}\kappa_{0}^{-{\overline{c}}},}$
${{}|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}F_{3}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|+\frac{1}{\widetilde{\rho}}|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}F_{3}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|\leq{\overline{\gamma}}\varepsilon_{0}\Big{(}\frac{\widetilde{\rho}}{\rho_{0}}\Big{)}^{2}\kappa_{0}^{-{\overline{c}}}}$
for ${\overline{\gamma}},{\overline{c}}$ constants (dimensionless), $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})\in{\mathcal{C}}(\widetilde{\rho},\kappa_{0})$.
[…]
Thus the corresponding $F^{\prime}_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}){\buildrel def\over{=}}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ admit the same bounds as Eq.2.18, but in the smaller domain ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ with $\delta_{0}\leq{\overline{\gamma}}(\frac{\varepsilon_{0}}{C})\kappa_{0}^{-{\overline{c}}}<\frac{1}{2}\kappa_{0}$, by simply suitably increasing the constants ${\overline{\gamma}},{\overline{c}}$, as discussed in AppendixB .
[…]
Given $\rho_{0},\kappa_{0},n\geq 1,\lambda=\frac{1}{2}$ define, for ${\overline{\gamma}},{\overline{c}}$, as in Eq.2.18:
$\rho_{n}=\Big{(}\frac{\varepsilon_{n-1}}{C}\Big{)}^{\lambda}\rho_{n-1},\quad\kappa_{n}=\kappa_{n-1}-2\delta_{n-1},\quad\delta_{n}={\overline{\gamma}}\frac{\varepsilon_{n}}{C}\kappa_{n}^{-{\overline{c}}}$ (2.19)
Iterate the renormalization transformation $({{\bf A}}_{n-1},{\mbox{\boldmath$\alpha$}}_{n-1})\to({{\bf A}}^{\prime}={{\bf A}}_{n},{\mbox{\boldmath$\alpha$}}^{\prime}={\mbox{\boldmath$\alpha$}}_{n})$ transforming the Hamiltonian $H_{n-1}({{\bf A}}_{n-1},{\mbox{\boldmath$\alpha$}}_{n-1})$, on ${\mathcal{C}}(\rho_{n-1},\kappa_{n-1})$ with pertubation $f_{n-1}$, into $H_{n}({{\bf A}}_{n},{\mbox{\boldmath$\alpha$}}_{n})$, on ${\mathcal{C}}(\rho_{n},\kappa_{n})$ with pertubation $f_{n}$. Suppose inductively, for all $n\geq 0$:
$\gamma_{0}\varepsilon_{n}J_{0,-}^{-1}<\rho_{n},\qquad C^{-1}\rho_{n}\ J_{0,+}<1,\qquad\kappa_{n}>\frac{1}{2}\kappa_{0}$ (2.20)
Then, for $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})\in{\mathcal{C}}(\rho_{n},\kappa_{n})$, the $\varepsilon_{n}=\|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}f_{n}\|_{\rho_{n},\kappa_{n}}+\frac{1}{\rho_{n}}\|{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\alpha$}}}f_{n}\|_{\rho_{n},\kappa_{n}}$ are bounded, if $\kappa_{n}>\frac{1}{2}\kappa_{0}$, by:
${{}\frac{\varepsilon_{n}}{C}=\gamma\Big{(}\frac{\varepsilon_{n-1}}{C}\Big{)}^{\xi}{(\frac{\kappa_{0}}{2})^{-c}},\qquad\kappa_{n}=\kappa_{n-1}-\delta_{n-1}}$
with $\gamma,c$ dimensionless, $\xi=\frac{3}{2}$.
The recursion, Eq. 2.19 , for $\frac{\varepsilon_{n}}{C}$ can be iterated for $n\geq 1$, as long as the conditions Eq.2.10 remain valid, finding $\frac{\varepsilon_{n}}{C}\leq\Lambda(\frac{\varepsilon_{n-1}}{C})^{\xi}$ for $\Lambda={\overline{\gamma}}(\frac{\kappa_{0}}{2})^{-{\overline{c}}}$ and $\gamma,c,\xi=\frac{3}{2}$ dimensionless constants. If $\eta_{n}{\buildrel def\over{=}}\sigma\frac{\varepsilon_{n}}{C}$ with $\sigma=\Lambda^{2}$ this becomes simply $\eta_{n}\leq\eta_{0}^{\xi^{n-1}}$, and consequently $\rho_{n}\leq\Lambda^{-\frac{n(n+1)}{2}}\eta_{0}^{\xi^{n}-1}\rho_{0}=O(\eta_{0}^{\xi^{n}})\rho_{0}=O((\Lambda^{2}\frac{\varepsilon_{0}}{C})^{\xi^{n}})\rho_{0}$.
Then Eq.2.19 imply that $\kappa_{n}=\kappa_{0}-2\sum_{k=1}^{n}{\overline{\gamma}}\frac{\varepsilon_{n-k}}{C}(\frac{1}{2}\kappa_{0})^{-{\overline{c}}}$ is bounded by $\kappa_{n}=\kappa_{0}-2\sum_{k=0}^{n-1}\Lambda^{-1}\eta_{k}$: therefore the conditions in Eq.2.20 can be reproduced if $\eta_{0}=\sigma\frac{\varepsilon_{0}}{C}=\Lambda^{2}\frac{\varepsilon_{0}}{C}$ is small enough.
The result is that in the coordinates $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}){\buildrel def\over{=}}({\bf A}_{n},{\mbox{\boldmath$\alpha$}}_{n})$, up to a time scale $\sim{\mbox{\boldmath$\omega$}}_{0}^{-1}O(\Lambda^{2}\Big{(}\frac{\varepsilon_{0}}{C}\Big{)}^{-\xi^{n}})$ and in region of phase space of order $\rho_{n}\sim\rho_{0}\Big{(}\Lambda^{2}\frac{\varepsilon_{0}}{C}\Big{)}^{\xi^{n}}$, the motion is quasi periodic motion with spectrum ${\mbox{\boldmath$\omega$}}_{0}$ and Hamiltonian $H_{n}$ with perturbation size $\sim\varepsilon_{n}\sim(\Lambda^{2}\frac{\varepsilon_{0}}{C})^{\xi^{n}}$.
The conclusion is that analyticity in the $\bf A$}} variables is lost: in the limit $n\to\infty$ remains a quasi periodic motion confined for time $\in(0,+\infty)$ on the real (and analytic) torus to which the approximations ${{\bf A}}^{\prime}_{n},{\mbox{\boldmath$\alpha$}}^{\prime}_{n}$ lead.
  • It is also possible to define a sequence of maps $\widetilde{\cal K}_{n}$ defined in the fixed domain ${\mathcal{C}}_{\frac{1}{4}\rho_{0},\frac{1}{4}\kappa_{0}}$ by rescaling the action coordinates of the polydisks by a factor $(\frac{\varepsilon_{n-1}}{C})^{-\frac{1}{2}}=\rho_{n-1}/\rho_{n}$, $n\geq 1$ so that they are all turned into ${\mathcal{C}}_{\frac{1}{4}\rho_{0},\frac{1}{4}\kappa_{0}}$: calling $\widetilde{\eta}_{n}{\buildrel def\over{=}}\frac{\varepsilon_{n-1}}{C}$ the rescaling transformation will rescale time by $\widetilde{\eta}_{n}^{-\frac{1}{2}}$ and will change ${\bf A}_{n}$ into ${\bf A}^{\prime}_{n}=\widetilde{\eta}_{n}^{-\frac{1}{2}}{\bf A}_{n}$ and change the Hamiltonian into:
    $\widetilde{H}_{n}={\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}^{\prime}_{n}+\widetilde{\eta}_{n}^{\frac{1}{2}}\frac{1}{2}({\bf A}^{\prime}_{n}\cdot J_{0}{\bf A}^{\prime}_{n})+\widetilde{\eta}_{n}^{-\frac{1}{2}}f_{n}(\widetilde{\eta}_{n}^{\frac{1}{2}}{\bf A}^{\prime}_{n},{\mbox{\boldmath$\alpha$}}^{\prime}_{n})\,\vtop{\halign{#\cr\rightarrowfill\crcr\kern-1.0pt\nointerlineskip\cr\;${\scriptstyle n\to\infty}$\;\crcr}}\,{\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}^{\prime}_{\infty}$ (2.22)
    and in the rescaled variables the sizes of the anharmonic terms tend to 0 superexponentially, taking into account the recursion defined in Eq.(2.20) which implies that the size of $f_{n}$ is of order $\frac{\varepsilon_{n}}{C}=(\widetilde{\eta}_{n})$.
  • Hence the perturbation $f_{0}$ and the twist $J_{0}$ are, after renormalization, “irrelevant” operators (in Eq.(2.17) both tend to 0 as $n\to\infty$), while the harmonic oscillator ${\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}$ is a “fixed point”: in some sense the transformation has the harmonic oscillator as an attractive fixed point. This completes a proof of the KAM theorem, interpreted in the Renormalization Group frame [4, 20, 21, 22]: it can be classified as a “super-renormalizable” problem, as it requires only a second order perturbation analysis around the trivial fixed point, Eq.(2.7).
Remarks:
(1) The analysis of this section follows closely [23] correcting a few dimensional typos.
It is possible to redefine $J_{1}\equiv J_{0}+\frac{1}{2}{\partial}^{2}_{\bf a}{\overline{f}}({\bf a})$ and correspondingly add to $F_{1}$, in Eq.(2.5), the second order term in $\bf A$$^{\prime}$ namely ${{\bf A}}^{\prime}\cdot\frac{1}{2}{\partial}^{2}_{\bf a}{\overline{f}}({\bf a}){{\bf A}}^{\prime}$. This modifies $F_{0}$ in the first line of Eq. 2.5 replacing $J_{0}$ with $J_{1}$ and the improves the bound of $F^{\prime}_{1}$ to: $|F_{1}^{\prime}({{\bf A}}^{\prime})|<\gamma_{1}\varepsilon_{0}(\frac{\widetilde{\rho}}{\rho_{0}})^{2}$. Then defining $\rho_{n}=(\frac{\varepsilon_{0}}{C})^{\lambda}\rho_{n-1}$ with $\lambda=\frac{1}{3}$ it follows, as above, $\frac{\varepsilon_{n}}{C}<\Lambda(\frac{\varepsilon_{0}}{C})^{\frac{4}{3}}$ (as, in this case, the best choice of $\lambda$ is $\lambda=\frac{1}{3}$ so that $\xi=\frac{4}{3}>\frac{3}{2}$).
In the ${{\bf A}}_{n},{\mbox{\boldmath$\alpha$}}_{n}$ the motion is controlled by a quadratic Hamiltonian ${\mbox{\boldmath$\omega$}}_{0}{{\bf A}}_{n}+\frac{1}{2}{{\bf A}}_{n}J_{n}{{\bf A}}_{n}$ in a portion of action space $O((\frac{\varepsilon_{n}}{C})^{\frac{1}{3}}\rho_{0})$ ”much larger” compared to the previous case, where the size was $O(\frac{\varepsilon_{0}}{C})^{\frac{1}{2}}\rho_{0}$, but $J_{n}=J_{0}$ was simpler.
The implication is that up a time scale of order $\omega_{0}^{-1}(\frac{\varepsilon_{0}}{C})^{-\frac{4}{3}n}$ the motion is quasi periodic in a domain of size $\rho_{0}(\frac{\varepsilon_{0}}{C})^{(\frac{1}{3})^{n}}$ controlled by Hamiltonian with quadratic part $J_{n}\neq J_{0}$. The previous choice gave control up to a time shorter $\omega_{0}^{-1}(\frac{\varepsilon_{0}}{C})^{-(\frac{3}{2})^{n}}$ controlled by the simpler Hamiltonian with quadratic part $J_{n}=J_{0}$ but in a domain of much smaller size $\rho_{0}(\frac{\varepsilon_{0}}{C})^{(\frac{1}{2})^{n}}$.
(2) The role of the constant $C$ is essential, of course, but it should be remarked that it appears in the smallness condition of $\frac{\varepsilon_{0}}{C}$ and in $J_{0,+}\rho_{0}C^{-1}<1$: both conditions can be fulfilled if $C=O(\varepsilon_{0}^{\xi})$ with $\xi<1$. Question: could this be of interest (aside the simple application in the remark in the following Sectiom, see  3.Sec.3).
(3) The analysis of the first iteration is extremely simplified if $f_{0}$ depends only on $\alpha$: this is a Thirring model and in Eq.2.5 it is ${\bf a}=0$, $F_{1}=0$, $F_{2}$ contains only the quadratic term but $F_{3}$ still has to be defined via Eq.2.8 so that at the next iteration dependence on $\bf A$$^{\prime}$ is introduced in $f_{1}$. Nevertheless the above proof remains valid and yields existence of the invariant torus with spectrum ${\mbox{\boldmath$\omega$}}_{0}$ without any condition on the lower bound $J_{0,-}$ as long as it is >0>0. The invariant tori in Thirring models are called “twistless” as no upper bound on $J_{0,-}$ is necessary: for such models a direct proof based on the explicit computation of a series, “Birkhoff series”, converging to the parametric equations of the invariant torus, can be derived, [24] , and covers also some cases in which $J_{0,-}=0$.  [^13] Which do not fit immediately in the above proof: the latter can be adapted to the case $H={\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}+\frac{1}{2}{{\bf A}}^{\perp}\cdot J_{0}{{\bf A}}^{\perp}+f_{0}({{\bf A}}^{\perp},{\mbox{\boldmath$\alpha$}})$ where ${{\bf A}}^{\perp}=(A_{1},A_{2},\ldots,A_{p},0,\ldots 0)$ if $J_{0,j}>0$ for $j\leq p$ and $(J_{0})_{j}=0$ for $j>p$, which generalizes the Thirring models and is even an easy extension of the twistless Hamiltonians theory (as the perturbation may now depend also on $\bf A$$^{\perp}$): the extension is also implied by the analysis in [24] .
3 FPUT
The FPUT digital work (reprinted in [25]), as soon as the contemporary KAM theorem became widely known, was qualitatively “explained”, at least by several physicists, [5], by the remark that the observed motions were generated by a “small” perturbation of the simple harmonic chain and could probably be treated via the results of the KAM theory: which, at the same time, was leading to a deeper understanding of the three body problem. The latter problem, as well as the FPUT chain, were in fact both perturbations of integrable systems, although affected by “resonances”.
However it took a long time to develop a complete proof that low energy motions observed in some FPUT experiments should mostly be quasi periodic, [26].
The models originally chosen in FPUT had Hamiltonian:
$H=\sum_{j=1}^{N-1}\frac{1}{2}p_{j}^{2}+W_{\alpha,\beta}(q_{j}-q_{j+1}),\qquad W_{\alpha,\beta}(x)=\frac{1}{2}x^{2}+\alpha x^{3}+\beta x^{4}$ (3.1)
with fixed endpoints $q_{0}=0,q_{N}=0$, [5]. If $\beta=0$ or $\alpha=0$ the models are respectively called fput-$\alpha$-chain or fput-$\beta$-chain.
The $\alpha$-chain, as a dynamical system, presents the difficulty that $W_{\alpha,0}(x)$, is unbounded below, if $\alpha>0$, or above if $\alpha<0$: hence it is not even clear whether all or, at least, some of the motions remain in a bounded phase space region (in a finite or infinite time). The problem does not arise in the $\beta$-chains if $\varepsilon>0$ as it is always supposed here.
In the original FPUT work nine choices of $\alpha$ and $\beta$ are considered and the cases numbered $4,5,6,7,9$ are apparently $\beta$ models (i.e. $\alpha=0,\beta>0$) while those numbered $1,2,3,8$ are $\alpha$ models ($\alpha>0,\beta=0$). In all simulations the motions kept a total energy constant, “within $1\%$ or so”.
The moving particles coordinates, N1{N}-1 in number, can be expressed as $q_{k}=\frac{\sqrt{2}}{\sqrt{N}}\sum_{h=1}^{N-1}\sin(\frac{\pi hk}{N})Q_{h}$ and $p_{k}=\frac{\sqrt{2}}{\sqrt{N}}\sum_{h=1}^{N-1}\sin(\frac{\pi hk}{N})P_{h}$ for $k=0,\ldots,N$.
The well known theory of the vibrating string with elastic potential energy $\sum_{k=0}^{N-1}\frac{1}{2}(q_{k}-q_{k+1})^{2}=\frac{1}{2}\sum_{h=1}^{N-1}\omega_{h}^{2}Q_{h}^{2}$ induces to study its oscillations via Forurier’s transforms coordinates $(\bf P,\bf Q)$ expressed as functions of $({\bf p},{\bf q})$ to which are canonically conjugated. If $\omega_{h}=2\sin\frac{\pi h}{N}$, therefore:
$H_{0}({\bf p},{\bf q})=H_{0}({\bf P,\bf Q})=\sum_{k=1}^{N-1}\frac{1}{2}(P_{k}^{2}+\omega_{k}^{2}Q_{k})^{2}\,{\buildrel def\over{=}}\,\sum_{k=1}^{N-1}\omega_{k}a_{k}$ (3.2)
where $a_{k}=\frac{(P_{k}^{2}+\omega_{k}^{2}Q_{k}^{2})}{2\omega_{k}}$. The latter is a natural coordinate as the pairs $(a_{k},\varphi_{k})=(a_{k},\arcsin(\frac{\omega_{k}Q_{k}}{\sqrt{2\omega_{k}a_{k}}}))$ are action-angle variables and evolve, in the elastic chain, as $(a_{k},\varphi_{k})\to(a_{k},\varphi_{k}+\omega_{k}t)$; and $({\bf p},{\bf q})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({\bf P,\bf Q})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({\bf a},{\mbox{\boldmath$\varphi$}})$ are symplectic maps.
Therefore $Q_{k}=(\frac{2a_{k}}{\omega_{k}})^{\frac{1}{2}}\sin\varphi_{k}$ and, as an example, the fput-$\beta$-chain Hamiltonian can be written, if $s_{h}=\sin\frac{\pi h}{N},c_{h}=\cos\frac{\pi h}{N}$:
${{}H_{\varepsilon}({\bf p},{\bf q})=\sum_{k=1}^{N-1}\omega_{k}\alpha_{k}+\varepsilon\sum_{{\bf h}}T_{\bf h}\prod_{i=1}^{4}Q_{h_{i}}={\mbox{\boldmath$\omega$}}_{k}\cdot{\bf a}+\varepsilon f_{0}({\bf a},{\mbox{\boldmath$\varphi$}})}$
${{}T_{\bf h}=(\frac{2}{N})^{2}\sum_{k=1}^{N-1}\prod_{j=1}^{4}(s_{h_{j}k}(1-c_{h_{j}})-c_{h_{j}k}s_{h_{j}}),}$
${{}f_{0}({\bf a},{\mbox{\boldmath$\varphi$}})=\frac{1}{2^{4}}\sum_{{\bf h}\in Z^{4},{\mbox{\boldmath$\sigma$}}}T_{\bf h}\prod_{j=1}^{4}\sigma_{j}e^{i\sigma_{j}\varphi_{h_{j}}}(\frac{2a_{h_{j}}}{\omega_{j}})^{\frac{1}{2}},}$
where ${\mbox{\boldmath$\sigma$}}=(\sigma_{j})_{j=1}^{4},\sigma_{j}=\pm 1$ and the variables $Q_{k}$ have been expressed in terms of the ${\bf a},{\mbox{\boldmath$\varphi$}}$.
The first mathematical results, [17], dealt with the $\beta$-chain with fixed extremes $q_{0}=q_{N}=0$, considered as a perturbation of order $\varepsilon$ of the elastic chain with N1N-1 particles (i.e. of the chain with $\alpha=0,\beta=0$).
Perturbation theory can be applied to $H_{\varepsilon}$: just look for a canonical transformation generated by a function $\varepsilon\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ as ${\bf a}={\bf a}^{\prime}+\varepsilon{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\varphi$}}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}),{\mbox{\boldmath$\varphi$}}^{\prime}={\mbox{\boldmath$\varphi$}}+\varepsilon{\mbox{\boldmath$\partial$}}_{{\bf a}^{\prime}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ which, if possible, transforms $H_{\varepsilon}({{\bf A}},{\mbox{\boldmath$\varphi$}})$ into $H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ (where $\varphi$ is imagimed expressed as function of $({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$).
$H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})={\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon{\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\varphi$}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}),{\mbox{\boldmath$\varphi$}})+\varepsilon f_{0}({\bf a}^{\prime}+\varepsilon{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\varphi$}}}\Phi,{\mbox{\boldmath$\varphi$}})$ (3.4)
Naturally $\Phi$ would be chosen so that $H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ is ${\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}$. Calculating up to $O(\varepsilon^{2})$ this means that $\Phi$ should be such that $\varepsilon{\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\varphi$}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})+\varepsilon f_{0}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ vanishes. However even this might be impossible if there are ${\mbox{\boldmath$\nu$}}\in Z^{N-1}$ such that ${\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\nu$}}=0$.@
Therefore the best that can be done is to subtract from the above $\Delta_{4}$ the terms which, expressed in the Fourier’s series, have “resonant” harmonics (orthogonal to $\omega$ but $\neq 0$) and define $\Phi$ to cancel, at lowest order, the part of $f_{0}$ left. So $\Phi$ can be chosen so that the new Hamiltonian (once laboriously expressed in the variables $({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$) will contain, up to order $\varepsilon^{2}$, only resonant harmonics (including ${\mbox{\boldmath$\nu$}}=0$).
The new Hamiltonian will have the form $H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})={\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon{\overline{H}}_{1}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})+\varepsilon^{2}H_{2}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$ where ${\overline{H}}_{1}$ is entirely resonant.
Repeating the argument $s$-times the Hamiltonian is represented by $H_{\varepsilon}=H_{2}+\sum_{k=1}^{s}\varepsilon^{s}{\overline{H}}_{s}+\varepsilon^{s+1}H_{\varepsilon,s+1}$ where the ${\overline{H}}_{s}$ contain only resonant terms, which remain the same if $s$ is increased. The series obtained in this way is the “Birkhoff series”, which however is generically divergent as shown by Poincaré, [27] .
However truncating the sum to an order $s\geq 1$ and neglecting the $\varepsilon^{s+1}H_{\varepsilon,s+1}$ might turn out to be a simpler system. This happens in the present case if $N$ is prime or is not a power of $2$, as pointed out in [17], simply because there are no ${\mbox{\boldmath$\nu$}}\neq 0$ with ${\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}=0$ in the Fourier’s expansion of the potential (i.e. no resonances $\nu$ with $|{\mbox{\boldmath$\nu$}}|\leq 4$). The terms of order $\leq\varepsilon^{2}$ in ${\overline{H}}_{1},{\overline{H}}_{2}$ not only do not depend at all on the angles $\mbox{\boldmath$\varphi$}$$^{\prime}$, but ${\overline{H}}_{1}=0$ and the terms in ${\overline{H}}_{2}$ can be collected in the form $\varepsilon^{2}{\bf a}^{\prime}\cdot J{\bf a}^{\prime}$, where $J$ is explicitly computed in terms of the spectrum ${\mbox{\boldmath$\omega$}}_{0}$ of the elastic chain and furthermore it is $\det J\neq 0$.
Hence the Hamiltonian ${\overline{H}}_{\varepsilon}={\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon^{2}{\overline{H}}_{2}({\bf a}^{\prime})$ is integrable and, as computed in [17], is found to be:
${\overline{H}}_{\varepsilon}({\bf a}^{\prime})=\sum_{k=1}^{N-1}\omega_{k}a^{\prime}_{k}+\frac{\varepsilon^{2}}{2N}\Big{(}\sum_{h,k}^{1,N-1}\frac{\omega_{h}\omega_{k}}{8}\ a^{\prime}_{h}a^{\prime}_{k}-\sum_{k}\frac{\omega_{k}^{2}}{32}{a^{\prime}_{k}}^{2}\Big{)}$ (3.5)
The surprising simplicity of this normal form, and the $\omega_{h}\neq 0,h=1,\ldots,N-1$, immediately allows to check that ${\mbox{\boldmath$\omega$}}\cdot{\bf a}+\frac{\varepsilon}{2}{\bf a}\cdot J{\bf a}$ has a Jacobian matrix $J$ constant and $\det J\neq 0$.
The few values of $N$, i.e. other than a prime or a power of $2$, which were still missing has been solved in [26] where a careful analysis shewed that no restriction should be set on the $N$ of fput-$\beta$-chains with fixed extremes, and identified the basic reason for the existence of an integrable Birkhoff transform permitting to avoid the issue of resonances affecting non trivially the Birkhoff transform, see also [28] .
The difficulty remaing after [17] was that, given $N$, it was not excluded that resonant harmonics were absent in the Fourier expansion of $H_{2}$: the only way out seemed to compute the corresponding Fourier’s coefficients. A task that could be performed (as it was done as time elapsed) also for single choices of $N$, but not covering all values of $N$.
The breakthrough idea, [18], is to forget (temporarily) the computation of the normal form and to check that the resonant harmonics, that could contribute resonances to the normal form, could not be present because they break symmetries of the Hamiltonian (which are preserved in the normal form construction). Thus the normal form can be constructed correctly as if the assumption in [17] could be valid (although it was known to be false in few interesting cases: for examples (with $N=15,21$) see [28, Sec.8.5] . The symmetries are briefly presented in AppendixC.
The spectra of the motions of ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ will have the form ${\mbox{\boldmath$\omega$}}={\mbox{\boldmath$\omega$}}_{0}+\varepsilon J{\bf a}^{\prime}$ and $KAM$ can be applied to perturbations of the considered integrable model. Let $\Omega$ be an open region in phase space, in which the ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ is defined: then $\Omega$ is covered by the tori corresponding to the actions $\bf a$}: e.g. $0<H_{\varepsilon}({\bf a}^{\prime})<E$.
The $\det J\neq 0$ implies that the set in $\Omega$ of invariant tori for whose spectra verify the inequality $|{\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\nu$}}|>C|{\mbox{\boldmath$\nu$}}|$ fails, for some $0\neq{\mbox{\boldmath$\nu$}}\in Z^{N}$, has volume $C\varepsilon^{-2}\kappa|\Omega|$, for a constant $\kappa>0$: hence the for almost points in $\Omega$ the spectrum $\omega$ verifies a Diophantine inequality.
Hence under a perturbation of ${\overline{H}}_{\varepsilon}$ of size $\eta$ a large part of phase space will remain covered by invariant tori quasi periodically run with $C_{\varepsilon}=C\varepsilon^{2}$-Diophantine frequencies if $C_{\varepsilon}$ is large enough compared to $\eta$. The $\beta$-chain being approximated to order $\eta=\varepsilon^{2}$ by the above integrable system ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ will therefore have a large part of a set $Q$ in phase space covered by $C_{\varepsilon}=C\varepsilon^{2}$-Diophantine tori. In particular this applies to the set $Q$ of energy close to the ground state of the linear chain where $H({\bf p},{\bf q})<E$ (if $E>0$ and $\varepsilon$ is small enough), and the fraction of $Q$ covered by invariant tori will approach $1$ as $\varepsilon\to 0$ at fixed $E>0$. Remark: (1) The statement that KAM implies that at small energy $H_{\varepsilon}({\bf p},{\bf q})<E$ “most phase space points are on invariant tori” requires some care. Describing the system in the coordinates $({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$ the set ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ depends on $\varepsilon$: but the application of KAM to ${\overline{H}}\varepsilon({\bf a}^{\prime})$ imposes a condition on the same $\varepsilon$. The remark at the end of Sec.2 is useful here, as it stresses that if the KAM theory is applied to cases in which the Diophantine constant has the form $C\varepsilon^{-1}$ the KAM theorem holds provided the perturbation size $\eta\varepsilon^{-1}$ with $h$ small enough.
(2) About the $\alpha$-model it is possible that, applying perturbation theory to the fput-$\alpha$-chain, an equivalent Hamiltonian is obtained which is an $O(\varepsilon^{2})$ perturbation of an integrable Hamiltonian with $J_{\varepsilon}$ still of $O(\varepsilon)$: but this is not (yet?) proved, [28]. See also the heuristic analysis in Appendix D.
(3) The approach to the $\beta$-model at fixed extremes can be applied to consider the Toda lattice chains with periodic or fixed boundary conditions as approximations to the fput-$\beta$-chains models. The above analysis, [26], suggests to try to use integrability of the Toda lattice to play the pertubation theory role in the search of an integrable Birkhoff approximation.
The suggestion is to consider the fput-Hamiltonians as perturbations of the Toda-lattice and consequences of integrability of several forms of its Hamiltonian:
${{}H({\bf p},{\bf x})=\sum_{j=1}^{N-1}\frac{1}{2}p_{j}^{2}+\sum_{j=0}^{N-1}W_{T}(q_{j}-q_{j+1})}$
${{}W_{T}(x)=\frac{1}{\varepsilon^{2}}e^{\varepsilon x}=\frac{1}{\varepsilon^{2}}(1+\varepsilon x)+\frac{1}{2!}x^{2}+\frac{\varepsilon}{3!}x^{3}+\frac{\varepsilon^{2}}{4!}x^{4}+\frac{\varepsilon^{3}}{5!}x^{5}+\ldots}$
integrable for instance with boundary conditions $q_{0}=q_{N}$, periodic lattice, [29, 30, 31], or $q_{0}=-\infty,q_{N}=+\infty$, lattice with endpoints fixed at $\pm\infty$, [32], or with fixed endpoints $q_{0}=q_{N}=0$.
In a sense the idea is that the Toda chains might play the role of an integrable Birkhoff normal form, of which fput-chains can be considered a perturbation. In AppendixD the basic results, [31], on the normal form of the periodic Toda-chains are summarized and heuristically applied to the fput-chains.
Appendices
Appendix A Shift equation
Existence of the shift ${\bf a}=-J_{0}^{-1}{\partial}_{\bf a}{\overline{f}}_{0}({\bf a})$.
Remark the bound on the matrix elements $\sigma_{ij}=J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{{\bf a}}^{2}{\overline{f}}_{0}({\bf a})$ in $C_{\lambda\rho_{0}}$ by $\sigma=((1-\lambda)\rho_{0})^{-1}J_{0,-}^{-1}\varepsilon_{0}$: hence the map $S{\bf a}={\bf a}+J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}({\bf a})$ has non singular Jacobian if $N\sigma<\frac{1}{3}$ (for instance)  141414 Let $|v|=\sum_{i}|v_{i}|$ for $v\in C^{N}$. Then the matrix $\delta_{ij}+\sigma_{ij}$ applied to $v\in C^{N}$ results in $w_{i}=v_{i}+\sum_{j}\sigma_{ij}v_{j}$: i.e. $|w|=\sum_{i}|w_{i}|\geq|v|(1-N\sigma)>0$ if, say, $N\sigma<\frac{1}{3}$.  and the map is locally invertible; furthermore any pair ${\bf a}^{\prime},{\bf a}^{\prime\prime}\in C((1-\lambda)\rho_{0})$ is mapped into a pair $S{\bf a}^{\prime},S{\bf a}^{\prime\prime}$ at distance $|S{\bf a}^{\prime}-S{\bf a}^{\prime\prime}|>(1-\sigma N)|{\bf a}^{\prime}-{\bf a}^{\prime\prime}|$, showing invertibility of $S$ in $C_{(1-\lambda)\rho_{0}}$ with image in $C((1+N\sigma)(1-\lambda)\rho_{0})\subset C(\rho_{0})$ if $\lambda$ is fixed $\lambda=\frac{2}{3}$ and $N\sigma<\frac{1}{3}$.
As $\bf a$} varies on the boundary of $C_{(1-\lambda)\rho_{0}}$ the image $S{\bf a}$ describes a set surrounding the polydisk of radius $(1-N\sigma)(1-\lambda)\rho_{0}=\frac{2}{9}\rho_{0}$: therefore the images $S{\bf a}$ cover a set containing 0 and imply existence of $\bf a$} with $S{\bf a}=0$ and therefore, $|{\bf a}|=|J_{0}^{-1}{\partial}_{\bf a}{\overline{f}}_{0}({\bf a})|<J_{0,-}^{-1}\varepsilon_{0}$. In particular if $\lambda=\frac{2}{3},\gamma_{0}=3N$:
$3NJ_{0,-}^{-1}\varepsilon_{0}\rho_{0}^{-1}<1\Rightarrow|{\bf a}|<J_{0,-}^{-1}\varepsilon_{0}<\frac{1}{\gamma_{0}}\rho_{0},\qquad{\bf a}\in C(\frac{1}{3N}{\rho_{0}})$ (a.1)
in general any other $\lambda<1$ would be acceptable provided $\gamma_{0}=N(1-\lambda)^{-1}$; the choices of $\lambda,\gamma_{0}$ are just convenient for later use and far from optimal.
Appendix B Renormalized H: dimensional estimates
Remark that in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0})$ the Jacobian of the map ${\mbox{\boldmath$\alpha$}}^{\prime}={\mbox{\boldmath$\alpha$}}+{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}$ is $(I+{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0})$ and, from the bound on $\Phi_{0}$, Eq.2.12 , $|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0}|<6\widehat{g}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}$.   [^15] Recall that $\Phi_{0}$ is bounded in ${\mathcal{C}}(\frac{2}{3}\rho_{0},\kappa_{0})$ and $|{\bf a}|<\frac{1}{6}\rho_{0}$; then use $\frac{2}{3}-\frac{1}{2}=\frac{1}{6}$. By Eq.2.12, $|{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})|<\widehat{\gamma}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-\widehat{c}}$ and the Jacobian determinant will be $\geq 1-\delta_{0}>0$ if $\delta_{0}\leq N!6\widehat{g}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}$ is smaller than $1$: for definitness suppose
$\delta_{0}=N!6\widehat{g}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}<\frac{1}{2}\kappa_{0}<1.$ (b.1)
Hence, for such $\delta_{0}$, the ${\mbox{\boldmath$\alpha$}}^{\prime}=S{\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}+{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, maps (by Eq. 2.12) , ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-\delta_{0})$ into a set $\subset{\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0})$ covering the polydisk ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ and is also locally invertible (as its Jacobian is nonsingular) and globally one-to-one, hence holomorphic.  161616 Remark that, for the purpose of solving the implicit functions problems in Eq.2.4 , an argument can be used similar to the one for the solution of ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ in Appendix A . Any pair ${\mbox{\boldmath$\alpha$}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime\prime}\in C(\frac{1}{2}\kappa_{0})$ is mapped, if $|v|=\sum_{i}|v_{i}|,v\in C^{N}$, into a pair $S{\mbox{\boldmath$\alpha$}}^{\prime},S{\mbox{\boldmath$\alpha$}}^{\prime\prime}$ at distance $|S{\mbox{\boldmath$\alpha$}}^{\prime\prime}-S{\mbox{\boldmath$\alpha$}}^{\prime}|>(1-N\|{\partial}_{{\mbox{\boldmath$\alpha$}}}{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}\|_{\frac{1}{2}\rho_{0},\kappa_{0}})\,|{\mbox{\boldmath$\alpha$}}^{\prime}-{\mbox{\boldmath$\alpha$}}^{\prime\prime}|$; hence the map $S$ is one-to-one globally and analyticity of $S^{-1}$ will also follow from the positivity of the Jacobian. For a more general and systematic study of the inversion of the maps in Eq.2.4 see Appendicx G in [33]. $S^{-1}$ will be denoted ${\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ and $\Delta$ inherits the bound Eq.2.13 $\|\Delta\|_{\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0}}<{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-{\overline{c}}}$ provided the conditions in Eq.(2.10) hold together with $\delta_{0}<\frac{1}{2}\kappa_{0}$, and the constants ${\overline{\gamma}},{\overline{c}}$ are chosen large enough. : its inverse in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ will be written ${\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ and fullfills the bound  2.13  for suitable ${\overline{\gamma}},{\overline{c}},\delta_{0}$.
For the purpose of dimensional bounds it is convenient to consider the second map $S$ in Eq.(2.4) in the domain ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ with $\delta_{0}$ in Eq.b.1 and values in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-\delta_{0})$.
The conditions for the bounds are summarized as, Eq. 2.10b.1:
$\gamma_{0}\varepsilon_{0}J_{0,-}^{-1}<\rho_{0},\qquad C^{-1}\rho_{0}\ J_{0,+}<1,\qquad\delta_{0}=\widetilde{\gamma}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-\widetilde{c}}<\frac{1}{2}\kappa_{0}$ (b.2)
for $\gamma_{0},\widetilde{\gamma},\widetilde{c}$ suitable dimensionless constants.  [^17] Likewise the ${{\bf A}}={{\bf A}}^{\prime}+{\bf a}+{\partial}_{\mbox{\boldmath$\alpha$}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ would lead to ${{\bf A}}={{\bf A}}^{\prime}-{\bf a}+\Xi({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ with $\Xi({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=-{\partial}_{\mbox{\boldmath$\alpha$}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))$ under similar conditions; but $\Xi$ is not needed here.
Given a generic function $F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ define $F^{\prime}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}){\buildrel def\over{=}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))$. Then, by Eq.2.20, to estimate the derivatives ${\partial}_{{{\bf A}}^{\prime}}$ and ${\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}$ of $F_{j}$ consider first the ${\partial}_{{{\bf A}}^{\prime}}$ and ${\partial}_{{\mbox{\boldmath$\alpha$}}}$. Then make use of, see also Eq. 2.102.13 :
${{\partial}_{{{\bf A}}^{\prime}}F^{\prime}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=}$ ${{}{\partial}_{{{\bf A}}^{\prime}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})+{\partial}_{\mbox{\boldmath$\alpha$}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\cdot{\partial}_{{{\bf A}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})}$
${{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}F^{\prime}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=}$ ${{}{\partial}_{{\mbox{\boldmath$\alpha$}}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\cdot(I+{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))}$
${\|\Delta\|_{\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0}}<}$ ${{}{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}}}$
${|{\partial}_{{{\bf A}}^{\prime}}\Delta|_{\frac{1}{4}\rho_{0},\kappa_{0}-2\delta_{0}}}$ ${{}+\frac{4}{\rho_{0}}|{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}\Delta|_{\frac{1}{4}\rho_{0},\kappa_{0}-2\delta_{0}}<{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}}}$
Hence $f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\equiv\sum_{j=1}^{3}F^{\prime}_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ and Eq.2.19:
${{}|{\partial}_{{{\bf A}}^{\prime}}f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})|=|{\partial}_{{{\bf A}}^{\prime}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})+{\partial}_{{\mbox{\boldmath$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}){\partial}_{{{\bf A}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})|}$
${{}\leq{\overline{\gamma}}\varepsilon_{0}(\frac{\varepsilon_{0}}{C})^{\frac{1}{2}}\kappa_{0}^{-{\overline{c}}}+\widetilde{\gamma}\rho_{0}\varepsilon_{0}(\frac{\varepsilon_{0}}{C})^{\frac{1}{2}}\kappa_{0}^{-{\overline{c}}}\frac{1}{\rho_{0}}{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}}}$
${{}|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})|=|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{\boldmath$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})(I+{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))|}$
${{}\leq|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{\boldmath$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})|(1+N{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}})}$
Hence the $\varepsilon_{1}$ can be bounded as the $\varepsilon^{\prime}_{1}$ in Eq.2.20 in the domain ${\mathcal{C}}(\frac{1}{4}\rho_{0},\kappa_{0}-4\delta_{0})$ (smaller than ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ to allow using dimensional bounds in the bounds of the derivatives of $\Delta$) simply by replacing the constants $\gamma_{j},c_{j}$ by a single pair ${\overline{\gamma}},{\overline{c}}$.
Appendix C FPUT symmetry
The Hamiltonian $H_{0}({\bf p},{\bf q})=\sum_{k=1}^{N}\frac{1}{2}(p_{k}^{2}+(q_{k}-q_{k+1})^{2})$ for $N=2(n+1)$ particles can be solved by changing variables $^{\bf q}\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,\widehat{\bf q},{\bf p}\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,\widehat{\bf p}$ with:
$q_{h}=\frac{1}{\sqrt{N}}\sum_{k=1}^{N}e^{\frac{2\pi i}{N}hk}\widehat{q}_{k},\quad p_{h}=\frac{1}{\sqrt{N}}\sum_{k=1}^{N}e^{\frac{2\pi i}{N}hk}\widehat{p}_{k}$ (c.1)
with $\widehat{q}_{h}={\overline{\widehat{q}}}_{N-h}$ and $\widehat{p}_{h}={\overline{\widehat{p}}}_{N-h}$ and periodic boundary $\widehat{q}_{N}=\widehat{q}_{0},\widehat{p}_{0}=\widehat{p}_{N}$.
It becomes $H_{0}(\widehat{\bf p},\widehat{\bf q})=\sum_{k=1}^{N}\omega_{k}a_{k}$ where $\omega_{k}=2\sin\frac{\pi k}{N}$ and $a_{k}=\frac{1}{2}(|\widehat{p}_{k}|^{2}+\omega_{k}^{2}|\widehat{q}_{k}^{2}|)$. The coordinates $\widehat{q}_{N},\widehat{p}_{N}$ describe the free motion of the center of mass; discarding $\widehat{q}_{N},\widehat{p}_{N}$ the Hamiltonian is reduced to 2n=N22n=N-2 degrees of freedom.
It is convenient to separate real and imaginary part of the $\widehat{\bf q}$ and $\widehat{\bf p}$ calling $Q_{h}=-\widehat{q}_{h,im}$ for $h\leq\frac{N}{2}$, and $Q_{\frac{N}{2}+h}=\widehat{q}_{h,re}$. Then
$q_{k}=\frac{\sqrt{2}}{\sqrt{N}}\sum_{h=1}^{N/2-1}(Q_{h}\sin\frac{2\pi kh}{N}+Q_{\frac{1}{2}N+h}\cos\frac{2\pi kh}{N})$ (c.2)
and define similarly $P_{h}$ in terms of the $\widehat{\bf p}$: in this way the phase space coordinates for $H_{0}$ will be denoted ${\bf Q}=(Q_{1},\ldots,Q_{\frac{1}{2}N},Q_{\frac{1}{2}N+1}\ldots,Q_{N})=({\bf Q_{+}},{\bf Q_{-}})$ and, likewise $\bf P=({\bf P_{+}},{\bf P_{-}})$.
Define the canonical maps $S,R$:
[…]
where ${\bf q}^{t}=(q_{2},\ldots,q_{N-1},q_{1})$ if ${\bf q}=(q_{1},\ldots,q_{N-1})$ and likewise is defined ${\bf p}^{t}$; the map $S$ is written for the variables $\bf Q$ to make clear that $S{\bf Q}={\bf Q}$ means, by Eq(c.2), that the manifold $F$ in phase space of the fixed points of the linear symplectic map $S$ consists of the ${\bf Q_{-}}=0,{\bf P_{-}}=0$ (i.e. of the ${\bf p},{\bf q}$ with purely imaginary Fourier’s tranforms $\widehat{\bf p},\widehat{\bf q}$). And from Eqc.2 it is seen that on $F$ it is $q_{N}=q_{0}=0$: furthermore the maps $R,S$ commute with $H_{\varepsilon}$ therefore evolution generated by $H_{\varepsilon}$ leaves invariant the manifold $E$.
Hence the $N$-degrees of freedom periodic b.c. Hamiltonian $H_{\varepsilon}$ describes, if restricted to $F$, a free chain with fixed endpoints and $n$ degrees of freedom if $N=2(n+1)$. Furthermore its Hamiltonian can be written:
$H_{0}=\sum_{k=1}^{n}\omega_{k}a_{k},\quad a_{k}=\frac{1}{2}\frac{(P_{k}^{2}+\omega_{k}^{2}Q_{k}^{2})}{\omega_{k}},\quad\sin\varphi_{k}=\frac{\omega_{k}Q_{k}}{\sqrt{2\omega_{k}a_{k}}}$ (c.4)
with $(a_{k},\varphi_{k})_{k=1}^{n}$ pairs of conjugate symplectic coordinates $({\bf a},{\mbox{\boldmath$\varphi$}})$.
The relation between the periodic chain, with $N$ particles, and the fixed extremes chain, with $n=\frac{N}{2}-1$ particles, leads to determine classes of integers $N$ for which $N$-particles chains Birkhoff’s series of some low order are integrable at small $\varepsilon$, [26], including various cases beyond those with NN=power of $2$ or $N$-prime.
Soon, [18], an original approach was developed to cover all chains with $N$ arbitrary integer (which by the time had still remained to relatively few integers): and the result was achieved without the need to compute the normal form, but using the symmetries, to infer in general, that fput-$\beta$-chains normal forms could not be affected by possible resonances that, altough possible, were conjectured not to be involved in the normal form construction.
It is interesting to remark that the the potential of fput-$\beta$-chain with fixed extremes, setting ${\mbox{\boldmath$\sigma$}}=(\sigma_{i})_{i=1}^{4},\sigma_{i}=\pm 1$ and $T_{\bf h}=(\frac{2}{N})^{2}\sum_{k=1}^{\frac{N}{2}}\prod_{j=1}^{4}((1-c_{h_{i}})s_{h_{i}k}-s_{h_{i}}c_{h_{i}k})$ for ${\bf h}=(h_{i})_{i=1}^{4}$, where $c_{h}=\cos\varphi_{h},s_{h}=\sin\varphi_{h}$. can be written:
$V_{4}({\bf Q})=\frac{1}{4}\sum_{{\bf h},{\mbox{\boldmath$\sigma$}}}T_{\bf h}\prod_{j=1}^{4}\sigma_{j}e^{i\sigma_{j}\varphi_{h_{j}}}(\frac{{a_{h_{j}}}}{{\omega_{h_{j}}}})^{\frac{1}{2}}$ (c.5)
and easily identifies the terms which need to be absent to obtain a normal form which is integrable: if $n_{{\bf h},{\mbox{\boldmath$\sigma$}}}(k)=\sum_{h_{j}=k}\sigma_{j}$ and ${\bf n}=(n_{{\bf h},{\mbox{\boldmath$\sigma$}}}(k))_{k=1}^{N=1}$ the uneliminable resonances $\nu$ are the ones for which $\xi={\bf n}\cdot{\mbox{\boldmath$\omega$}}=0$. Hence to prove the result it is necessary to show that the sum of the terms labeled by the above $\bf n$} for which $\xi=0$ either do not arise (but they do in interesting cases) or that they add up to 0 as a consequence of the symmetry or because ${\bf n}=0$, as found in [18, 28].
Appendix D Toda and FPUT: heuristics
The last suggestion in Sec.3 is to try to take advantage of the very detailed studies on the Toda chain described below. For this purpose a periodic Toda Hamiltonian will be written, [31]:
$H({\bf p},{\bf q})=\sum_{k=1}^{N}\Big{(}\frac{1}{2}p_{k}^{2}+\alpha^{2}e^{q_{k}-q_{k+1}}\Big{)}=\sum_{k=1}^{N}(\frac{1}{2}p_{k}^{2}+\alpha^{2}V_{T}(q_{k}-q_{k+1}))$ (d.1)
with $q_{1}=q_{N+1}$. This is obtained by a canonical transformation on Eq.3.2: $p=\varepsilon p^{\prime}$, $q=\frac{1}{\varepsilon}q^{\prime}$ followed by a rescaling of time $t^{\prime}=\varepsilon^{2}t$ leading to Eq.3.4, where $\alpha$ is a new coupling equal to $\varepsilon^{-2}$.
The periodic Toda lattice model is studied in [31] where $2N$ global canonical action-angle variables $({\bf a},{\mbox{\boldmath$\varphi$}})\in(R_{+}^{N},[0,2\pi]^{N})$ are constructed and evolve at time $t$ into $({\bf a},{\mbox{\boldmath$\varphi$}}+{\mbox{\boldmath$\omega$}}({\bf a})t)$, following a Hamiltonian $H_{\alpha}({\bf a})$. Near ${\bf a}=0$ it is $\omega_{i}({\bf a})=2\alpha\omega_{0,i}+\frac{1}{4N}a_{i}+O(|{\bf a}|^{2})$, if ${\mbox{\boldmath$\omega$}}_{0}$ is the spectrum of the linear oscillations $\omega_{0,i}=2\sin\frac{\pi i}{N}$.
The $H_{\alpha}({\bf a})$, [31, th.1.4], is globally defined for ${\bf a}\in R^{N}_{+}$ and the Jacobian $J_{\alpha}({\bf a})=\det\frac{{\partial}{\mbox{\boldmath$\omega$}}({\bf a})}{{\partial}{\bf a}}$ is analytic in $\alpha>0$; for $\alpha>0$ it is strictly convex in ${\bf a}\in R_{+}^{N}$ and $J_{\alpha}({\bf a})_{ij}=\frac{1}{4N}\delta_{ij}+O(|{\bf a}|)\neq 0$ and the spectrum is $\omega_{i}=2\alpha\sin(\frac{i\pi}{N})+\frac{1}{4N}a_{i}+O(|{\bf a}|^{2})$.
The general KAM theory of Sec.2 can be applied, as in Sec.3, to perturbations of the Toda-model because the Jacobian $J_{\alpha}$ strict convexity for $\alpha$ in any interval bounded away from 0 implies that the spectrum has the Dipohantine property: so pertubations of size $\eta$ will leave most of the invariant tori only slightly deformed if $\eta$ is small.
However application to the fput-chains requires more analysis mainly because the use of the change of variables leads to the Hamiltonian in Eq.c.1 with $\alpha$ small and compare it with Eq.3.6 $\varepsilon$ small: aside from the time rescaling, this means to compare motions with $\alpha$ in Eq.c.1 to motions with $\varepsilon$ in Eq.3.6 setting $\alpha=\varepsilon^{-2}$; so when $\alpha$ is very large.
The analysis in [31] is global and the coordinates $({\bf a},{\mbox{\boldmath$\varphi$}})$ exist for all $\alpha>0$ and transform analytically the Hamiltonian Eq.c.1 into a function $\widetilde{H}_{\alpha}({\bf a})$ with convex Jacobian $J_{ij}=\frac{{\partial}^{2}_{a_{i}}H({\bf a})}{{\partial}^{2}a_{j}}$, see th.(1.1),th.(1.2) in [31]. Hence in the region of couplings relevant for the fput-chains, i.e. $\varepsilon$ close to 0, $\alpha=\varepsilon^{-2}$ large, the Toda lattice Hamiltonian admits action-angles coordinates.
Proceeding heuristically: the Toda lattice can be considered a perturbation of order $\varepsilon^{2}$ of the fput-$\alpha$-lattice, see Eq.3.6:
$H_{3}({\bf p},{\bf q})=\sum_{j=0}^{N-1}\Big{(}\frac{1}{2}(p_{k}^{2}+(q_{k}-q_{k+1})^{2})+\frac{1}{3!}\varepsilon(q_{k}-q_{k+1})^{3}\Big{)}$ (d.2)
Hence the elastic chain perturbed by $h_{3}=\frac{1}{3!}\varepsilon\sum_{j=0}^{N-1}(q_{k}-q_{k+1})^{3}$ and the Toda lattice truncated to order $\varepsilon^{2}$ are the same: thus their first order Birkhoff normal forms, to remove the resonant terms present in $h_{3}$, coincide up to the next order $\varepsilon^{2}$ (the first step in the construction of Birkhoff normal form only involves the common part $h_{3}$). The integrability of the Toda lattice should then imply that the normal form for the above truncation is integrable implying that the normal form of the fput-$\alpha$-chain is integrable.
Therefore the above argument of [26] can be applied (heuristically here, and avoiding any explict calculation of the Birkhoff normal form) to reach the same conclusion and the periodic fput-$\alpha$-chain model, and consequently (again by the symmetry in [26]) the fixed extremes fput-$\beta$-chain, show quasi periodic motions, filling large regions of phase space in any region $\Omega$ where $e<H<E$, up to a set of volume approching 0 as $\varepsilon\to 0$.
The $\alpha$-chain can be studied in the same way, still heuristically: the periodic fput-chain with potential $V_{3,4}(x)=\frac{\varepsilon}{3!}x^{3}+\frac{\varepsilon^{2}}{4!}x^{4}$ is the same as the Toda lattice truncated at second order. Then the Birkhoff’s normal form of second order should be the same as that of the Toda lattice: therefore should be integrable. However the second order normal form of the $V_{3,4}$ fput-chain contains no cross terms between the two contributions to the potential (which have at least order $3$). Hence the two contributions to the mormal form of $V_{3,4}$ fput-chain are the same as those for the simple $V_{3}$-chain ($\alpha$-chain) to second order plus the first order ones of the $V_{4}$ (the $\beta$ chain which is already known to be integrable); hence the integrability of the Toda lattice allowing to express its normal form in terms of the action variables implies the same properties for the Birkhoff’s normal forms of the $V_{3,4}$ and $V_{3}$ fput chains.
The argument can be extended to chains with potential $V(x)=\sum_{k=3}^{s}c_{s}x^{s}$ for $s>4$ but only if $c_{s}=\frac{1}{s!}$, while for $s=3,4$ the $c_{3},c_{4}$ could be quite arbitrary.